Search the web.. GADroplet Of Ink In An × + | O Living in Japan: Lite Of An l choquestionspdf Course Home MastenngPhysics CH08 Home work-Microsoft Edge emView?assignmentProblemiD 87771681offset prev CH 08 Homework Think and Solve 8.44 e previous Think and Solve 8.44 Part A A droplet of ink in an industrial ink-jet printer carries a charge of 2.1 x 10-10 C and is deflected onto paper by a force of 4.0× 10-4 N Find the strength of the electric field (E = F/q) produce this force Express your answer to two significant figures a appropriate units. EValue Units Submit My Answers Give Up Spider Man Adobe Photoshop Elements 2018 Suicide Squad Extended Cut Star Walk 2 Live Home 3D Pro Planes, Tra Automobi Blood Wars... H9 15999 1499 5A9 3999 519.99
Answer #1

electric field, E = F/q

= 4*10^-4 / 2.1*10^-10

= 1.9 * 106 N/C